OpenStudy (anonymous):

Solving Quadratics can you check if this is right if not what will it be? thanks 1) y=x^2+6x+8 y=(x+3)^2 -1 y=(x+2)(x+4) y=0 when x=-2 and x=-6 Has a minimum y value of -4 2) y=x^2-8x+12 y=(x+4)^2 -4 y=(x-2)(x-6) y=0 when x=2 and x=6 Has a minimum y value of -1

jhonyy9 (jhonyy9):

in case of first x why equal -6 ?

OpenStudy (anonymous):

positive +6

jhonyy9 (jhonyy9):

not not will be never 6

OpenStudy (anonymous):

Yeah, see basically factorize. 1. the first one can be factorized by splitting the middle term. the second one, imagine x + 3 as 1 entitiy. so you can resolve it just like x^2 - 1 is resolved. so the second eq reduces to (x+3+1)(x+3-1)

OpenStudy (anonymous):

If you can factorize, the answer comes out like a piece of cake

OpenStudy (anonymous):

Are they in correct form so far?

jhonyy9 (jhonyy9):

so than y=(x+2)(x+4) how you solve now ?

OpenStudy (anonymous):

x^2+6x+8

jhonyy9 (jhonyy9):

how you solve after have got it factored completly like (x+2)(x+4) ?

OpenStudy (solomonzelman):

\(\large\color{black}{ 1)~~~y=x^2+6x+8 }\) \(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{ ~~~~~~~y=x^2+6x+8+1-1 }\) \(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{ ~~~~~~~y=x^2+6x+9-1 }\) \(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{ ~~~~~~~y=(x+3)^2-1 }\) \(\LARGE\color{white}{ \rm \left| \right| }\) the minimum value is at: \(\large\color{black}{ ~(-3,-1) }\) \(\LARGE\color{white}{ \rm \left| \right| }\)

jhonyy9 (jhonyy9):

so first x+2=0 x=-2 and x+4 =0 so x=-4 ok ?

OpenStudy (solomonzelman):

yes, the x-intercepts are at (-2,0) and (-4,0) as it is shown above.

OpenStudy (solomonzelman):

can you find the minimum value and the x-intercepts correctly for the second problem?

OpenStudy (solomonzelman):

x-intercepts in number 2 are correct.

OpenStudy (solomonzelman):

\(\large\color{black}{ 2)~~~y=x^2-8x+12 }\) \(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{ ~~~~~~~y=x^2-8x+12+4-4 }\) \(\LARGE\color{white}{ \rm \left| \right| }\) \(\large\color{black}{ ~~~~~~~y=x^2-8x+16-4 }\) \(\LARGE\color{white}{ \rm \left| \right| }\)

OpenStudy (solomonzelman):

can you find the vertex?

OpenStudy (anonymous):

So are they right? i wasnt sure what the minimum of the y value was for each

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

How do you know if it has a minimum y value of -4 or -1?